OpenStudy (anonymous):

How do I solve this problem? Solve for cosθ in sin^2θ + cos^2θ = 1

terenzreignz (terenzreignz):

This? \[\large \sin^2\theta+\cos^2\theta=1\]

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Well, let's see... How to break this to you... cosθ can literally be any value from [-1 , 1] Let me explain... the value of cosθ depends on the value of θ, right?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Well, unfortunately (fortunately?) ANY value of θ would make this statement true sin²θ + cos²θ = 1

hartnn (hartnn):

he/she needs to solve for cos so, subtract sin^2 term from both sides, what u get ?

terenzreignz (terenzreignz):

Oops... Must have been overexcited :) My bad

OpenStudy (anonymous):

Im trying to write it but its giving me a hard time

OpenStudy (anonymous):

±√(1 + sin2Θ)

hartnn (hartnn):

did you mean \(\pm \sqrt{1-\sin^2 \theta}\) ??

OpenStudy (anonymous):

yes I pasted the wrong one Sorry, my computer is acting really crazy today

hartnn (hartnn):

ohh..if you meant that, then you are correct.

OpenStudy (anonymous):

ok thanks you made this very easy, thank you.. :)

hartnn (hartnn):

glad to hear :) welcome ^_^